OpenStudy (anonymous):

What is the kinetic energy of a 1.0-kg billiard ball that moves at 5.0 m/s? A. 5.0 J B. 6.5 J C. 10.0 J D. 12.5 J

OpenStudy (anonymous):

@kym02

OpenStudy (anonymous):

Do you remember the formula?

OpenStudy (anonymous):

i thiink so 1/2 times mass times (velocity squared) 12mv2

OpenStudy (anonymous):

he equation for finding Kinetic Energy KE is half the mass m times velocity v squared. KE=12mv2

OpenStudy (anonymous):

correct! so since\[K _{E}=\frac{ 1 }{ 2 }mv ^{2}\] then, \[\frac{ 1 }{ 2 } \times 1.0kg \times 5.0 ms ^{-1}\] \[=12.5 J\] So, the answer is D.

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Anytime!